# O. Ragniscoa, F.Zullob

### v1

@@ Abstract
In the present paper we derive a further extension of the results contained in two recent articles, both published in Open Communications in Mathematical Physics, where it was shown that the integrable version of the NN-species Volterra model, introduced by V. Volterra in 1937, is in fact maximally superintegrable. Here we point out that the superintegrability property applies as well to the case of infinitely many competing species, either countable or uncountable. Analytical and numerical results are given.
Keywords: Volterra integrable model for infinitely many species, Hamiltonian formulation, Superintegrable systems.
@@ 1 Introduction
The interest on dynamical systems of Lotka-Volterra type has been constantly growing all over the last century, starting from the pioneering results contained in the original articles [13, 14, 24, 25]. Among the extremely rich literature on the subject (see for instance [7], [23], [16], [3], [12], [4], [8],[22], [9], [5], [19], [6]), one paper deserves in our opinion a very special attention. It is the long essay published by Vito Volterra in 1937 [25], where some results of capital importance have been achieved: first, the original predator-prey model has been extended to an arbitrary number of species; second, the NN-species model has been framed within the context of Lagrangian and Hamiltonian mechanics; finally, a completely integrable subcase has been identified. Relying on these outstanding results, the authors of the present paper performed some further hopefully non-trivial steps: namely, they discovered that the integrable version introduced in [25] was in fact maximally superintegrable [17], [18], and as such reducible to a Hamiltonian system with only one degree of freedom. Here we make yet a further step, investigating the extension to infinitely many competing species, either countable or uncountable, always staying within the integrable Hamiltonian framework, and discover that the maximal superintegrability property is preserved by such generalization. Various different cases of the interaction operator are discussed, and a few periodic numerical examples are displayed. We may thus assert that Infinity collapses to One.
To this aim, in Section 2 we briefly recall the results obtained in [17], [18] for the integrable Volterra model in the case of a finite number of species. The new developments are contained in Sections 3 and 4. In particular, in Section 3 we consider the case of a denumerable collection of species, and in Section 4 we grapple with the problem of uncountable many species. Both Sections consist of 3 subsections, where we give first a general outlook and then face two different cases, concerning respectively even and odd kernel of the interaction operators. In section 5 we first present and analyze some analytical examples and then discuss numerical implementations. In Section 6 we critically analyze the results obtained and point out some affordable ways to improve the description of the system.
@@ 2 The case of a finite number N of species
The equations for the integrable $N$-species Volterra system have been introduced by V.Volterra in [25], and thoroughly investigated in [17], [18]:
$\frac{dN_{r}}{dt}=\epsilon_{r}N_{r}+\sum_{s\neq r=1}^{N}A_{rs}N_{r}N_{s}\quad\quad r=1,\ldots,N.$ (2.1)
In the equation (2.1), $\epsilon_{r}$ are the natural growth coefficients of each species and $A_{rs}$ are interaction coefficients between species $r$ and species $s$ that account for the effects of encountering between two individuals. In the integrable $N$-species case, the $N\times N$ interaction matrix $A$ has the form [18]:
$A_{rs}=\epsilon_{r}\epsilon_{s}(B_{r}-B_{s}),$ (2.2)
@@
$(z,\epsilon)=0,~(z,\eta)=1.$ (2.4)
It is convenient to rewrite the evolution equations (2.1) In terms of the logarithmic variables defined by $y_{j}:=\log N_{j}$, getting:
$\dot{y}_{j}=\epsilon_{j}+\sum_{l=1}^{N}A_{jl}\exp(y_{l})=\epsilon_{j}+(\eta_{j}\sum_{l=1}^{N}\epsilon_{l}-\epsilon_{j}\sum_{l=1}^{N}\eta_{l})\exp(y_{l})$ (2.5)
@@
$y_{n}=P\epsilon_{n}+Q\eta_{n}+k_{n},$ (2.6)
where in (2.6) $k_{n}$ is a generic element $k\in{\mathcal{K}}$. [^1] To be precise, with the term $k_{n}$ we mean $\sum_{k=1}^{N-2}R_{k}\tau_{m}^{(k)}$, i.e. the decomposition of $k_{n}$ along a given orthonormal basis. We adopt here this shorter symbol for easiness of notation. The evolution equations (2.5) take the form:
$\dot{P}\epsilon_{n}+\dot{Q}\eta_{n}+\dot{k}_{n}=\epsilon_{n}+\eta_{n}\sum_{k=1}^{N}\epsilon_{k}\exp(y_{k})-\epsilon_{n}\sum_{=1}^{N}\eta_{k}\exp(y_{k})$ (2.7)
@@
$\dot{P}=1-\sum_{n=1}^{N}\eta_{n}\exp(P\epsilon_{n}+Q\eta_{n}+k_{n}),$ (2.8)
$\dot{Q}=\sum_{n=1}^{N}\epsilon_{n}\exp(P\epsilon_{n}+Q\eta_{n}+k_{n}),$ (2.9)
$\dot{k}_{n}=0,\quad n=1...N-2$ (2.10)
@@
$\dot{P}=-\frac{\partial{\mathcal{H}}}{\partial Q};~~\dot{Q}=\frac{\partial{\mathcal{H}}}{\partial P}$ (2.11)
@@ 3.1 General outlook
In the discrete infinite case we extend the involved summation to the whole discrete line, taking care of convergence properties. Here we have to point out that, on one hand, we can look at the operator $iA$ as a self-adjoint separable Hilbert-Schmidt operator [11] mapping the Hilbert space $\ell^{2}$ of square summable sequences into itself, thanks to the the inner product $(x,y)\doteq\sum_{n\in{\mathbb{Z}}}x_{n}y_{n}$. As such, its spectrum consists of two opposite real eigenvalues, say $\pm\omega$, and a null eigenvalue of infinite multiplicity, the Hilbert space $\ell^{2}$ being spanned by the two orthonormal eigenvectors associated to $\pm\omega$ and by the orthogonal subspace Ker(A)(A) of codimension 2. On the other hand, $A$ acts nonlinearly on the space of sequences, according to the law $\dot{y}=A\exp(y)$, mapping $\ell^{\infty}$ into $\ell^{2}$. To be on the safe side, from now on we will assume that $\epsilon$ belongs to $\ell^{2}$, while $\eta$ and $\kappa$ belong to the dense open subset $S^{\infty}\subset\ell^{2}$ consisting of bounded sequences $s_{n}$ that vanish at $\infty$ faster than any power of $1/n$, i.e. such that $\lim_{n\to\infty}n^{p}s_{n}=0,\forall p$; this condition is quite natural for $\eta$, whose elements are defined by $\eta_{n}=B_{n}\epsilon_{n}$, and simply amounts to require that $B_{n}$ itself shares the same properties; its extension to $\kappa$ ensures that the whole quantity $Q\eta_{n}+\kappa_{n}$ enjoys the same regularity properties. Anyway, the possibility for $\eta$ and $\kappa$ to belong to other spaces will be also discussed in the next Sections.
@@
$\dot{y}_{n}=\epsilon_{n}+\sum_{m\in{\mathbb{Z}}}(\epsilon_{m}\eta_{n}-\eta_{m}\epsilon_{n})\exp(y_{m}).$ (3.1)
The sequence $y_{n}$ can be projected on the two complementary subspaces ${\mathcal{I}}=ImA$ and ${\mathcal{K}}=KerA$ in the form:
$y_{n}=P\epsilon_{n}+Q\eta_{n}+\kappa_{n}$ (3.2)
@@
$\sum_{n\in{\mathbb{Z}}}k_{n}\epsilon_{n}=0,\quad\sum_{n\in{\mathbb{Z}}}\kappa_{n}\eta_{n}=0.$ (3.3)
Since $\epsilon_{n},\eta_{n}$ and $\kappa_{n}$ both converge to zero as $|n|\to\infty$, the numerosity $N_{n}\to 1$ as $|n|\to\infty$. The equations (3.1) split in the following $3$ equations:
$\dot{P}=1-\sum_{m\in{\mathbb{Z}}}\eta_{m}\exp(P\epsilon_{m}+Q\eta_{m}+\kappa_{m})$ (3.4)
$\dot{Q}=\sum_{m\in{\mathbb{Z}}}\epsilon_{m}\exp(P\epsilon_{m}+Q\eta_{m}+\kappa_{m})$ (3.5)
$\dot{\kappa}_{n}=0$ (3.6)
@@
$0=1-\sum_{m\in{\mathbb{Z}}}\eta_{m}\exp(P^{0}\epsilon_{m}+Q^{0}\eta_{m}+\kappa_{m})$ (3.7)
$0=\sum_{m\in{\mathbb{Z}}}\epsilon_{m}\exp(P^{0}\epsilon_{m}+Q^{0}\eta_{m}+\kappa_{m})$ (3.8)
@@
Should we follow the alternative way of the spectral decomposition of the operator $A$ and apply the Hilbert-Schmidt machinery, as we did in [17] and [18] for the case of finite species, we had to assume that both the sequences $\epsilon_{n}$ and $\eta_{n}$ belong to $\ell^{2}$. We would have to look at the roots of the secular equation:
$\lambda_{\pm}v^{\pm}(n)=\alpha^{\pm}\eta_{n}-\beta^{\pm}\epsilon_{n}$ (3.9)
As a matter of fact, the nonzero eigenvalues read:
$\lambda_{\pm}\doteq\pm i\omega=\pm i{\sqrt{|\epsilon|^{2}|\eta|^{2}-(\epsilon,\eta)^{2}}}$ (3.10)
@@
$|\eta|^{2}\alpha-(\eta,\epsilon)\beta=\lambda\beta$ (3.11)
$(\epsilon,\eta)\alpha-|\epsilon|^{2}\beta=\lambda\alpha$ (3.12)
By solving for $\alpha,\beta$ the system (3.11) and (3.12) in terms of variables $\epsilon,\eta$ we get for the eigenfunctions $v_{n}^{\pm}\doteq u_{n}\pm iw_{n}$ the following expressions:
$u_{n}=\frac{\epsilon_{n}}{|\epsilon|},\quad w_{n}=\frac{1}{|\epsilon|\omega}\left((\epsilon,\eta)\epsilon_{n}-|\epsilon|^{2}\eta_{n}\right)$ (3.13)
@@
$A_{n,m}=i\omega(w_{m}u_{n}-w_{n}u_{m}).$ (3.14)
The projection of $y_{n}$, equivalent to (3.2), now reads
$y_{n}=pu_{n}-qw_{n}+\kappa_{n},$ (3.15)
whereas the equations of motion (3.1) reduce to
$\dot{p}=|\epsilon|+\omega\sum_{m\in\mathbb{Z}}w_{m}\exp(pu_{m}-qw_{m}+\kappa_{m}),$ (3.16)
$\dot{q}=-\omega\sum_{m\in\mathbb{Z}}u_{m}\exp(pu_{m}-qw_{m}+\kappa_{m}),$ (3.17)
$\dot{k}_{m}=0,$ (3.18)
@@
$\mathcal{H}=\omega\left(-|\epsilon|q+\sum_{m\in\mathbb{Z}}\exp(pu_{m}-qw_{m}+\kappa_{m})\right)$ (3.19)
In the next Section we will show how it is possible to give some very general expressions for the equations of motion in the infinite discrete case. This is somewhat surprising since in the case of finite species we have not been able to achieve such general results. By looking at equations (3.3), in order to be operative it is natural to assume that the sequences $\epsilon_{n}$ and $\eta_{n}$ have an opposite parity with respect to the Kernel sequence $\kappa_{n}$. So, in the following two subsections, we will consider the case of a Kernel given by an even sequence (and hence $\eta_{n}$ and $\epsilon_{n}$ odd sequences), and the case of a Kernel given by an odd sequence (and hence $\eta_{n}$ and $\epsilon_{n}$ even sequences). The conditions to get the convergence will be specified in the corresponding subsections.
@@ 3.2 The case of an even Kernel
$\begin{split}&\epsilon_{n}=\frac{\tilde{\epsilon}_{n}}{\sqrt{2N}},\quad|\tilde{\epsilon}_{n}|^{2}=1\quad\tilde{\epsilon}_{-n}=-\tilde{\epsilon}_{n},\\ &\eta_{n}=-\eta_{-n},\\ &\kappa_{n}=\kappa_{-n}\end{split}$ (3.20)
In the previous it is assumed that $\tilde{\epsilon}_{n}$ has compact support in [N,N][-N,N], so it is an odd function of $n$ assuming the values $\pm 1$ on the interval [N,N][-N,N]. The sequence $\eta_{n}$ is odd also, whereas $\kappa_{n}$ is even since it must be orthogonal to the sequences $\epsilon_{n}$ and $\eta_{n}$. Notice that $\epsilon_{n}$ is normalized in $\ell^{2}$ to 1, and the same normalization can be chosen for $\eta_{n}$. As previously mentioned, both $\eta_{n}$ and $\kappa_{n}$ are assumed to belong to the dense open subset $S^{\infty}$ of $\ell^{2}$ consisting in rapidly decreasing sequences. We start with the equation for $\dot{Q}$:
$\dot{Q}=\sum_{n\in{\mathbb{Z}}}\epsilon_{n}\exp(P\epsilon_{n}+Q\eta_{n}+\kappa_{n}).$ (3.21)
In order to deal with equation (3.21), we make use of the following
Proposition 3.1
@@
$S(P,x,N)=\sum_{n=-N}^{N}\epsilon_{n}\exp(P\epsilon_{n})f(xs_{n}),$ (3.22)
where $f(x)$ is a continuous function of $x$ and $\epsilon_{n}$ is given in (3.20). Then, for any finite values of $x$ one has
$\lim_{N\to\infty}S(P,x,N)=Pf(0).$ (3.23)
Proof: by taking the Taylor expansion of the exponential we get
$S(P,x,N)=\sum_{n=-N}^{N}\epsilon_{n}(1+P\epsilon_{n}+\frac{P^{2}}{2}\epsilon_{n}^{2}+\ldots)f(xs_{n}).$ (3.24)
@@
$\lim_{N\to\infty}P\sum_{n=-N}^{N}\epsilon_{n}^{2}f(xs_{n})=\lim_{N\to\infty}P\frac{1}{2N}\sum_{n=-N}^{N}f(xs_{n})=Pf(0).$ (3.26)
It follows that all the powers of $\epsilon$ greater than $2$ does not give a contribution in the Taylor series expansion of $\exp(P\epsilon_{n})$ in (3.22).
@@
$\sum_{n=-N}^{N}\epsilon_{n}f(xs_{n})=\frac{1}{\sqrt{2N}}\sum_{n=1}^{N}\tilde{\epsilon}_{n}\left(f(xs_{n})-f(xs_{-n})\right)$ (3.27)
Notice that, since $s_{n}$ is a sequence in $S^{\infty}$ $f(xs_{n})$ is a bounded sequence and in the limit $n\to\infty$ (and hence $N\to\infty$) it converges to $f(0)$ (the sequence $f(xs_{n})-f(0)$ is bounded and approaches zero). The above sum, by adding and subtracting $f(0)$, can be rewritten as
$\sum_{n=-N}^{N}\epsilon_{n}f(xs_{n})=\sum_{n=1}^{N}\epsilon_{n}\left(f(xs_{n})-f(0)\right)-\sum_{n=1}^{N}\epsilon_{n}\left(f(xs_{-n})-f(0)\right)$ (3.28)
@@
$|\frac{1}{\sqrt{2N}}\sum_{n=1}^{N}\tilde{\epsilon}_{n}\left(f(xs_{n})-f(0)\right)|\leq\frac{C|x|}{\sqrt{2N}}\sum_{n=1}^{N}|s_{n}|$ (3.29)
and, since $s_{n}\in S^{\infty}$ the sum is finite and in the limit $N\to\infty$ the linear term in $\epsilon_{n}$ in the Taylor series expansion of $\exp(P\epsilon_{n})$ in (3.22) vanishes. $\square$
@@
Corollary 3.2
The equation of motion (3.21) reduces, in the limit $N\to\infty$ to
$\dot{Q}=P.$ (3.30)
@@
$\dot{P}=1-\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(P\epsilon_{n}+Q\eta_{n}+\kappa_{n}).$ (3.31)
@@
$T(P,x,N)=\sum_{n=-N}^{N}\eta_{n}\exp(P\epsilon_{n})f(xs_{n}),$ (3.32)
where $f(x)$ is a continuous function of $x$ and $\epsilon_{n}$ is given in (3.20). Then, for any finite values of $x$ one has
$\lim_{N\to\infty}T(P,x,N)=\sum_{n\in\mathbb{Z}}\eta_{n}f(xs_{n})$ (3.33)
Proof: Indeed, if one consider the sum $\sum_{n=-N}^{N}\eta_{n}f(xs_{n})$, it is convergent in the limit $N\to\infty$. For large $n$ the terms $s_{n}\to 0$ so $f(xs_{n})\to f(0)$. Being continuous, $f(x)$ is locally bounded around $x=0$ so we can find an $M$ such that $|f(xs_{n})|<M$ for all values of $n$. It follows the convergence of the sum $\sum_{n=-N}^{N}\eta_{n}f(xs_{n})$ from the $S^{\infty}$ convergence of $\eta_{n}$. If one considers the Taylor expansion of $\exp(P\epsilon_{n})$ in the sum (3.32), only the constant term survives in the limit $N\to\infty$ and we get (3.33). $\square$
By using Proposition (3.3) we get for the expression (3.31)
$\dot{P}=1-\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}).$ (3.34)

### v2

# Open Communications in Nonlinear Mathematical Physics ]ocnmp[ Vol.6 (2026) pp 1References Article

@@ Abstract
In the present paper we derive a further extension of the results contained in two recent articles, both published in Open Communications in Nonlinear Mathematical Physics, where it was shown that the integrable version of the NN-species Volterra model, introduced by V. Volterra in 1937, is in fact maximally superintegrable. Here we point out that the superintegrability property applies as well to the case of infinitely many competing species, either countable or uncountable. Analytical and numerical results are given.
@@ 1 Introduction
The interest on dynamical systems of Lotka-Volterra type has been constantly growing all over the last century, starting from the pioneering results contained in the original articles [12, 13, 23, 24]. Among the extremely rich literature on the subject (see for instance [2, 4, 3, 5, 6, 7, 22, 8, 15, 11, 21, 18]), one paper deserves in our opinion a very special attention. It is the long essay published by Vito Volterra in 1937 [24], where some results of capital importance have been achieved: first, the original predator-prey model has been extended to an arbitrary number of species; second, the NN-species model has been framed within the context of Lagrangian and Hamiltonian mechanics; finally, a completely integrable subcase has been identified. Relying on these outstanding results, the authors of the present paper performed some further hopefully non-trivial steps: namely, they discovered that the integrable version introduced in [24] was in fact maximally superintegrable [16, 17], and as such reducible to a Hamiltonian system with only one degree of freedom. Here we make yet a further step, investigating the extension to infinitely many competing species, either countable or uncountable, always staying within the integrable Hamiltonian framework, and discover that the maximal superintegrability property is preserved by such generalization. Various different cases of the interaction operator are discussed, and a few periodic numerical examples are displayed. We may thus assert that Infinity collapses to One. Let us anticipate that, with a suitable choice of the interaction parameters, the one dimensional system becomes the classical one body Hamiltonian $\mathcal{H}=\frac{P^{2}}{2}+V(Q)$.
To this aim, in Section 2 we briefly recall the results obtained in [16], [17] for the integrable Volterra model in the case of a finite number of species. The new developments are contained in Sections 3 and 4. In particular, in Section 3 we consider the case of a denumerable collection of species, and in Section 4 we grapple with the problem of uncountable many species. Both Sections consist of different subsections, where we give first a general outlook and then face two different cases, concerning respectively even and odd kernel of the interaction operators. In section 5 we first present and analyze some analytical examples and then discuss numerical implementations. In Section 6 we recapitulate and critically analyze the results obtained and point out some affordable ways to improve the description of the system.
@@ 2 The case of a finite number N of species
The equations for the integrable $N$-species Volterra system have been introduced by V.Volterra in [24], and thoroughly investigated in [16, 17]:
$\frac{dN_{r}}{dt}=\epsilon_{r}N_{r}+\sum_{s\neq r=1}^{N}A_{rs}N_{r}N_{s}\quad\quad r=1,\ldots,N.$ (1)
In the equation (1), $N_{r}$ is the numerosity of the $r^{th}$ species, $\epsilon_{r}$ are the natural growth coefficients of each species and $A_{rs}$ are interaction coefficients between species $r$ and species $s$ that account for the effects of encountering between two individuals. In the integrable $N$-species case, the $N\times N$ interaction matrix $A$ has the form [17]:
$A_{rs}=\epsilon_{r}\epsilon_{s}(B_{r}-B_{s}),$ (2)
@@
$(z,\epsilon)=0,~(z,\eta)=1.$ (4)
It is convenient to rewrite the evolution equations (1) In terms of the logarithmic variables defined by $y_{j}:=\log N_{j}$, getting:
$\dot{y}_{j}=\epsilon_{j}+\sum_{l=1}^{N}A_{jl}\exp(y_{l})=\epsilon_{j}+(\eta_{j}\sum_{l=1}^{N}\epsilon_{l}-\epsilon_{j}\sum_{l=1}^{N}\eta_{l})\exp(y_{l})$ (5)
@@
$y_{n}=P\epsilon_{n}+Q\eta_{n}+k_{n},$ (6)
where in (6) $k_{n}$ is a generic element $k\in{\mathcal{K}}$. [^1] To be precise, with the term $k_{n}$ we mean $\sum_{k=1}^{N-2}R_{k}\tau_{m}^{(k)}$, i.e. the decomposition of $k_{n}$ along a given orthonormal basis. We adopt here this shorter symbol for easiness of notation. The evolution equations (5) take the form:
$\dot{P}\epsilon_{n}+\dot{Q}\eta_{n}+\dot{k}_{n}=\epsilon_{n}+\eta_{n}\sum_{k=1}^{N}\epsilon_{k}\exp(y_{k})-\epsilon_{n}\sum_{=1}^{N}\eta_{k}\exp(y_{k})$ (7)
@@
$\dot{P}=1-\sum_{n=1}^{N}\eta_{n}\exp(P\epsilon_{n}+Q\eta_{n}+k_{n}),$ (8)
$\dot{Q}=\sum_{n=1}^{N}\epsilon_{n}\exp(P\epsilon_{n}+Q\eta_{n}+k_{n}),$ (9)
$\dot{k}_{n}=0,\quad n=1...N-2$ (10)
@@
$\dot{P}=-\frac{\partial{\mathcal{H}}}{\partial Q};~~\dot{Q}=\frac{\partial{\mathcal{H}}}{\partial P}$ (11)
@@ 3.1 General outlook
In the discrete infinite case we extend the involved summation to the whole discrete line, taking care of convergence properties. Here we have to point out that, on one hand, we can look at the operator $iA$ as a self-adjoint separable Hilbert-Schmidt operator [10] mapping the Hilbert space $\ell^{2}$ of square summable sequences into itself, thanks to the the inner product $(x,y)\doteq\sum_{n\in{\mathbb{Z}}}x_{n}y_{n}$. As such, its spectrum consists of two opposite real eigenvalues, say $\pm\omega$, and a null eigenvalue of infinite multiplicity, the Hilbert space $\ell^{2}$ being spanned by the two orthonormal eigenvectors associated to $\pm\omega$ and by the orthogonal subspace Ker(A)(A) of codimension 2. On the other hand, $A$ acts nonlinearly on the space of sequences, according to the law $\dot{y}=A\exp(y)$, mapping $\ell^{\infty}$ into $\ell^{2}$. To be on the safe side, from now on we will assume that $\epsilon$ belongs to $\ell^{2}$, while $\eta$ and $\kappa$ belong to the dense open subset $S^{\infty}\subset\ell^{2}$ of rapidly decreasing sequences $s_{n}$ that vanish at $\infty$ faster than any power of $1/n$, i.e. such that $\lim_{n\to\infty}n^{p}s_{n}=0,\forall p$; this condition is quite natural for $\eta$, whose elements are defined by $\eta_{n}=B_{n}\epsilon_{n}$, and simply amounts to require that $B_{n}$ itself shares the same properties; its extension to $\kappa$ ensures that the whole quantity $Q\eta_{n}+\kappa_{n}$ enjoys the same regularity properties. Anyway, the possibility for $\eta$ and $\kappa$ to belong to other spaces will be also discussed in the next Sections.
@@
$\dot{y}_{n}=\epsilon_{n}+\sum_{m\in{\mathbb{Z}}}(\epsilon_{m}\eta_{n}-\eta_{m}\epsilon_{n})\exp(y_{m}).$ (13)
The sequence $y_{n}$ can be projected on the two complementary subspaces Im(A)(A) and Ker(A)(A) in the form:
$y_{n}=P\epsilon_{n}+Q\eta_{n}+\kappa_{n}$ (14)
@@
$\sum_{n\in{\mathbb{Z}}}k_{n}\epsilon_{n}=0,\quad\sum_{n\in{\mathbb{Z}}}\kappa_{n}\eta_{n}=0.$ (15)
Since $\epsilon_{n},\eta_{n}$ and $\kappa_{n}$ both converge to zero as $|n|\to\infty$, the numerosity $N_{n}\to 1$ as $|n|\to\infty$. The equations (13) split in the following $3$ equations:
$\dot{P}=1-\sum_{m\in{\mathbb{Z}}}\eta_{m}\exp(P\epsilon_{m}+Q\eta_{m}+\kappa_{m})$ (16)
$\dot{Q}=\sum_{m\in{\mathbb{Z}}}\epsilon_{m}\exp(P\epsilon_{m}+Q\eta_{m}+\kappa_{m})$ (17)
$\dot{\kappa}_{n}=0$ (18)
@@
$0=1-\sum_{m\in{\mathbb{Z}}}\eta_{m}\exp(P^{0}\epsilon_{m}+Q^{0}\eta_{m}+\kappa_{m})$ (19)
$0=\sum_{m\in{\mathbb{Z}}}\epsilon_{m}\exp(P^{0}\epsilon_{m}+Q^{0}\eta_{m}+\kappa_{m})$ (20)
@@
Should we follow the alternative way of the spectral decomposition of the operator $A$ and apply the Hilbert-Schmidt machinery, as we did in [16, 17] for the case of finite species, we had to assume that both the sequences $\epsilon_{n}$ and $\eta_{n}$ belong to $\ell^{2}$. We would have to look at the roots of the secular equation:
$\lambda_{\pm}v^{\pm}(n)=\alpha^{\pm}\eta_{n}-\beta^{\pm}\epsilon_{n}$ (21)
As a matter of fact, the non-zero eigenvalues read:
$\lambda_{\pm}\doteq\pm i\omega=\pm i{\sqrt{|\epsilon|^{2}|\eta|^{2}-(\epsilon,\eta)^{2}}}$ (22)
@@
$|\eta|^{2}\alpha-(\eta,\epsilon)\beta=\lambda\beta$ (23)
$(\epsilon,\eta)\alpha-|\epsilon|^{2}\beta=\lambda\alpha$ (24)
By solving for $\alpha,\beta$ the system (23) and (24) in terms of variables $\epsilon,\eta$ we get for the eigenfunctions $v_{n}^{\pm}\doteq u_{n}\pm iw_{n}$ the following expressions:
$u_{n}=\frac{\epsilon_{n}}{|\epsilon|},\quad w_{n}=\frac{1}{|\epsilon|\omega}\left((\epsilon,\eta)\epsilon_{n}-|\epsilon|^{2}\eta_{n}\right)$ (25)
@@
$A_{n,m}=i\omega(w_{m}u_{n}-w_{n}u_{m}).$ (26)
The projection of $y_{n}$, equivalent to (14), now reads
$y_{n}=pu_{n}-qw_{n}+\kappa_{n},$ (27)
whereas the equations of motion (13) reduce to
$\dot{p}=|\epsilon|+\omega\sum_{m\in\mathbb{Z}}w_{m}\exp(pu_{m}-qw_{m}+\kappa_{m}),$ (28)
$\dot{q}=-\omega\sum_{m\in\mathbb{Z}}u_{m}\exp(pu_{m}-qw_{m}+\kappa_{m}),$ (29)
$\dot{k}_{m}=0,$ (30)
@@
$\mathcal{H}=\omega\left(-|\epsilon|q+\sum_{m\in\mathbb{Z}}\exp(pu_{m}-qw_{m}+\kappa_{m})\right)$ (31)
In the next Section we will show how it is possible to give some very general expressions for the equations of motion in the infinite discrete case. This is somewhat surprising since in the case of finite species we have not been able to achieve such general results. In the infinite case we have been constrained to choose suitable parameters in order to achieve the convergence and this constraint led to the general expressions that we will give. In the Conclusions these aspects will be further commented. By looking at equations (15), in order to be operative it is natural to assume that the sequences $\epsilon_{n}$ and $\eta_{n}$ have an opposite parity with respect to the Kernel sequence $\kappa_{n}$. So, in the following two subsections, we will consider the case of a Kernel given by an even sequence (and hence $\eta_{n}$ and $\epsilon_{n}$ odd sequences), and the case of a Kernel given by an odd sequence (and hence $\eta_{n}$ and $\epsilon_{n}$ even sequences). From our point of view this is not restrictive: indeed we will show that it is possible to consider a mixed Hamiltonian flow that contains both the even functions and the odd ones. The conditions to get the convergence will be specified in the corresponding subsections.
@@ 3.2 The case of an even Kernel
$\begin{split}&\epsilon_{n}=\frac{\tilde{\epsilon}_{n}}{\sqrt{2N}},\quad|\tilde{\epsilon}_{n}|^{2}=1\quad\tilde{\epsilon}_{-n}=-\tilde{\epsilon}_{n},\\ &\eta_{n}=-\eta_{-n},\\ &\kappa_{n}=\kappa_{-n}\end{split}$ (32)
In the previous it is assumed that $\tilde{\epsilon}_{n}$ has compact support in [N,N][-N,N], so it is an odd function of $n$ assuming the values $\pm 1$ on the interval [N,N][-N,N]. The sequence $\eta_{n}$ is also odd, whereas $\kappa_{n}$ is even since it must be orthogonal to the sequences $\epsilon_{n}$ and $\eta_{n}$. Notice that $\epsilon_{n}$ is normalized in $\ell^{2}$ to 1, and the same normalization can be chosen for $\eta_{n}$. As previously mentioned, both $\eta_{n}$ and $\kappa_{n}$ are assumed to belong to the dense open subset $S^{\infty}$ of $\ell^{2}$ consisting in rapidly decreasing sequences. We start with the equation for $\dot{Q}$:
$\dot{Q}=\sum_{n\in{\mathbb{Z}}}\epsilon_{n}\exp(P\epsilon_{n}+Q\eta_{n}+\kappa_{n}).$ (33)
In order to deal with equation (33), we make use of the following
Proposition 3.1.
@@
$S(P,x,N)=\sum_{n=-N}^{N}\epsilon_{n}\exp(P\epsilon_{n})f(xs_{n}),$ (34)
where $f(x)$ is a continuous function of $x$ and $\epsilon_{n}$ is given in (32). Then, for any finite values of $x$ one has
$\lim_{N\to\infty}S(P,x,N)=Pf(0).$ (35)
Proof. By taking the Taylor expansion of the exponential we get
$S(P,x,N)=\sum_{n=-N}^{N}\epsilon_{n}(1+P\epsilon_{n}+\frac{P^{2}}{2}\epsilon_{n}^{2}+\ldots)f(xs_{n}).$ (36)
@@
$\lim_{N\to\infty}P\sum_{n=-N}^{N}\epsilon_{n}^{2}f(xs_{n})=\lim_{N\to\infty}P\frac{1}{2N}\sum_{n=-N}^{N}f(xs_{n})=Pf(0).$ (38)
It follows that all the powers of $\epsilon$ greater than $2$ do not give a contribution in the Taylor series expansion of $\exp(P\epsilon_{n})$ in (34).
@@
$\sum_{n=-N}^{N}\epsilon_{n}f(xs_{n})=\frac{1}{\sqrt{2N}}\sum_{n=1}^{N}\tilde{\epsilon}_{n}\left(f(xs_{n})-f(xs_{-n})\right)$ (39)
Notice that, since $s_{n}$ is a sequence in $S^{\infty}$, $f(xs_{n})$ is a bounded sequence and in the limit $n\to\infty$ (and hence $N\to\infty$) it converges to $f(0)$ (the sequence $f(xs_{n})-f(0)$ is bounded and approaches zero). The above sum, by adding and subtracting $f(0)$, can be rewritten as
$\sum_{n=-N}^{N}\epsilon_{n}f(xs_{n})=\sum_{n=1}^{N}\epsilon_{n}\left(f(xs_{n})-f(0)\right)-\sum_{n=1}^{N}\epsilon_{n}\left(f(xs_{-n})-f(0)\right)$ (40)
@@
$|\frac{1}{\sqrt{2N}}\sum_{n=1}^{N}\tilde{\epsilon}_{n}\left(f(xs_{n})-f(0)\right)|\leq\frac{C|x|}{\sqrt{2N}}\sum_{n=1}^{N}|s_{n}|$ (41)
and, since $s_{n}\in S^{\infty}$ the sum is finite and in the limit $N\to\infty$ the linear term in $\epsilon_{n}$ in the Taylor series expansion of $\exp(P\epsilon_{n})$ in (34) vanishes. $\square$
@@
Corollary 3.2.
The equation of motion (33) reduces, in the limit $N\to\infty$ to
$\dot{Q}=P.$ (42)
@@
$\dot{P}=1-\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(P\epsilon_{n}+Q\eta_{n}+\kappa_{n}).$ (43)
@@
$T(P,x,N)=\sum_{n=-N}^{N}\eta_{n}\exp(P\epsilon_{n})f(xs_{n}),$ (44)
where $f(x)$ is a continuous function of $x$ and $\epsilon_{n}$ is given in (32). Then, for any finite values of $x$ one has
$\lim_{N\to\infty}T(P,x,N)=\sum_{n\in\mathbb{Z}}\eta_{n}f(xs_{n})$ (45)
Proof. Indeed, if one consider the sum $\sum_{n=-N}^{N}\eta_{n}f(xs_{n})$, it is convergent in the limit $N\to\infty$. For large $n$ the terms $s_{n}\to 0$ so $f(xs_{n})\to f(0)$. Being continuous, $f(x)$ is locally bounded around $x=0$ so we can find an $M$ such that $|f(xs_{n})|<M$ for all values of $n$. It follows the convergence of the sum $\sum_{n=-N}^{N}\eta_{n}f(xs_{n})$ from the $S^{\infty}$ convergence of $\eta_{n}$. If one considers the Taylor expansion of $\exp(P\epsilon_{n})$ in the sum (44), only the constant term survives in the limit $N\to\infty$ and we get (45). $\square$
By using Proposition (3.3) we get for the expression (43)
$\dot{P}=1-\sum_{n\in{\mathbb{Z}}}\eta_{n}\exp(Q\eta_{n}+\kappa_{n}).$ (46)